Calculate the weight of a sample of impure oxalic acid that should be taken for titration by 0.5000 N NaOH so that the percentage of H2C2O4.2H2O will be twice the buret reading? 



1
Expert's answer
2021-07-07T05:02:05-0400

Molar Mass of NaOH = 39.997

= 0.5 × 39.997

= 19.999

= 19.999+60 = 79.999

Molar Mass of H2C2O4.2H2O = 126.07

= 126.07/79.999

= 1.576g